4. A soft-drink machine is regulated so that it discharges an average of 200 mls per cup. If the amount of drink is normally distributed with a standard deviation equal to 15 mls. Below what value do we get the smallest 30% of the cups? 


1
Expert's answer
2022-02-01T15:28:41-0500

Let X represent the amount of drink distributed.

"\mu = 200 \\

\sigma=15"

The value if we get the smallest 30% of the drinks

"P(X<x) = 0.3\\

P(Z< \frac{x-200}{15}) = 0.3 \\

P(Z< -0.68) = 0.3\\

\frac{x-200}{15} = -0.68 \\

x = -0.68 \times 15 + 200 \\

x=189.95"